OpenStudy (anonymous):

A bus travels from Oakridge to Hillsboro, a distance of 320 miles. The bus is traveling at 60 mph. The driver is required to contact the station in Hillsboro when he is between 50 and 100 miles from the station. During what time interval should he contact the station at Hillsboro

OpenStudy (anonymous):

t=d/s 50/60=5/6 100/60=5/3 beween 5/6h and 5/3h

OpenStudy (anonymous):

am not quite sure